OpenStudy (anonymous):

help please √64 over √3 ?

OpenStudy (anonymous):

what do you want done here?

OpenStudy (anonymous):

\[\frac{\sqrt{64}}{\sqrt{3}} = \frac{8}{\sqrt{3}} = \frac{8\sqrt{3}}{3}\]

OpenStudy (anonymous):

thanks